# Scrambled Halton Subsequences and Inverse Star-Discrepancy

### v1

@@ Abstract
Braaten and Weller discovered that the star-discrepancy of Halton sequences can be strongly reduced by scrambling them. In this paper, we apply a similar approach to those subsequences of Halton sequences which can be identified to have low-discrepancy by results from p-adic discrepancy theory. For given finite $N$, it turns out that the discrepancy of these sets is surprisingly low. By that known empiric bounds for the inverse star-discrepancy can be improved.
@@ 1 Introduction
i.e. the boxes in the supremum are not necessarily anchored at 00. It relates to the star-discrepancy via $D_{N}^{*}(y_{n})\leq D_{N}(y_{n})\leq 2D_{N}^{*}(y_{n})$. Thus, the two notions of discrepancy necessarily posses the same asymptotic behavior up to a factor.
@@
Recall that for an integer $b\geq 2$ the $b$-ary representation of $n\in{\mathbb{N}}$ is $n=\sum_{j=0}^{\infty}e_{j}b^{j}$ with $0\leq e_{j}=e_{j}(n)<b$. Based on the radical-inverse function, the van der Corput sequence in base $b$ is defined by $\varphi_{b}(n)=\sum_{j=0}^{\infty}e_{j}b^{-j-1}$ for all $n\in{\mathbb{N}}$. It is well-known, see e.g. [Nie92], that
$D_{N}^{*}(\varphi_{b}(n))\leq\frac{1}{N}+\frac{b+1}{2N}+\frac{b-1}{2\log(b)}\frac{\log(N)}{N}.$
This bound should be compared to (and is for small $b$ not too far off from) the record holder for the best known discrepancy as constructed in [Ost09] which satisfies
@@
In fact, this record holder is a generalized van der Corput sequence. In other words, van der Corput, their subsequences and generalizations may be regarded as prime candidates when looking for sequences with a particularly small star-discrepancy.
An easier-to-describe method than in [Ost09] how to (empirically) further reduce the discrepancy of van der Corput was introduced in [BW79]: for fixed $b\in\mathbb{N}$ choose an arbitrary permutation $\pi$ of $\{0,\ldots,b-1\}$ and define the scrambled van der Corput sequence by
@@
It is not difficult to prove the same bound for the star-discrepancy as for the standard van der Corput sequence but the additional parameter allows for (empiric) reduction of the discrepancy. In [BW79], concrete choices for $\pi$ were suggested for all primes $p\leq 53$.
The concept of scrambling may, of course, also be applied to subsequences of van der Corput and we obtain the following result which may be regarded as a generalization of Theorem 1.2 in [Wei24] in combination with Theorem 3 in [Mei68] as will become clear from its proof.
@@
Let $f(n)$ be a permutation polynomial $\bmod\ p^{2}$ for some prime number $p$, which means that $f(n)$ induces a bijection on $\mathbb{Z}/p^{2}\mathbb{Z}$, and let $\pi$ be a permutation of $\{0,\ldots,p-1\}$. Then the discrepancy of the sequence $(x_{n}):=(\varphi_{p,\pi}(f(n)))$ satisfies
@@
We call such sequences scrambled van der Corput subsequences. The easiest examples are of the form $f(n)=an+b$ with $\gcd(a,p)=1$. We call $a$ a shift and denote such a sequence by $\varphi_{p,\pi,a}$. Other possible choices for $f(n)$ (depending on $p$) may be found in [Wei24]. Although the theoretical bound does not guarantee a very small discrepancy, allowing the flexibility in choosing the parameters $\pi$ and $a$ in Theorem 1 reduces the star-discrepancy for given $N$ significantly as can be seen from Table 1 where our results are compared to the approach from [BW79]. The difference between the original van der Corput sequence and the scrambled one according to [BW79] is much larger (except for $p=2$), but moving to subsequences always but once reduced the star-discrepancy as well. In our experiments the gain exceeded $10\%$ most of the times.
@@
Table 1: Comparison of discrepancies for van der Corput like sequences for $N=100$ (left) and $N=1,000$ (right). For $\varphi_{p,\pi,a}(N)$ the smallest found values are listed when using $500$ different shifts, $500$ random permutations per shift.
For a fair comparison with the original and scrambled van der Corput sequence it should however be born in mind that the star-discrepancy was by our method minimized for a given $N$ which might mean that it is particularly small for the chosen $N$ but not for many other $n<N$. However, we observe in Figures 1 and 2 of the Appendix that both, the scrambled sequence as well as the scrambled subsequence, almost systematically outperform the original van der Corput sequence while the former two methods have the lowest star-discrepancy for approximately the same number of $n\leq 1,000$.
Finding sequences with a particularly small discrepancy is not only a relevant question in dimension $d=1$ but even of higher importance for $d>1$, where the star-discrepancy for a sequence $(x_{n})\subset[0,1]^{d}$ and the $d$-dimensional Lebesgue measure $\lambda_{d}(\cdot)$ is defined by
@@
where the supremum is taken over all $d$-dimensional intervals $B=[0,b_{1})\times\ldots\times[0,b_{d})$ with $0\leq b_{i}\leq 1$ for $i=1,\ldots,d$. It is widely conjectured that $D_{N}^{*}(x_{n})=O(\log(N)^{d}/N)$ is the best achievable order of convergence for the star-discrepancy but this conjecture has only been proven in the case $d=1$, see [Sch72]. Furthermore, we remind the reader that the star-discrepancy of sequences (infinitely many points) in dimension $d$ corresponds to that of point sets (finitely many points) in dimension $d+1$, see [Rot54]. The (extreme) discrepancy $D_{N}(x_{n})$ again allows arbitrary multi-dimensional half-open intervals in the supremum, i.e. without necessarily anchoring one base point at zero. As a generalization of the one-dimensional case, the inequality $D_{N}^{*}(x_{n})\leq D_{N}(x_{n})\leq 2^{d}D_{N}^{*}(x_{n})$ holds.
Sequences with a particularly small (star-)discrepancy are of great interest for high-dimensional integration tasks. Due to the Koksma-Hlawka inequality, see e.g. [KN74], the worst case error when approximating an integral by the average of some function evaluations depends linearly on the discrepancy of the evaluation points. This motivates why it makes sense to ask, what is the smallest star-discrepancy achievable for a given $N\in\mathbb{N}$ and we set
$D^{*}(N,d):=\inf\{D_{N}^{*}(P)\,:\,P\subset[0,1]^{d}\#P=N\}.$
or equivalently define the inverse star-discrepancy by
@@
which is the minimum number of sample points that guarantees a discrepancy bound of at most $\varepsilon>0$. Note that even small reductions of $N$ are of practical importance, if an expensive function $f$ is evaluated at these points to (numerically) calculate an integral. Alternatively decreasing the size of $N$ can be motivated by the numerical stability of certain regression problems which depend linearly on the discrepancy, see [WN19]. Therefore, precise theoretical bounds and numerical estimates of $N^{*}(\varepsilon,d)$ are of great interest.
As the asymptotic bound $\log(N)^{d}/N$ for the star-discrepancy exponentially depends on the dimension, it does not provide helpful information for moderate size sets. In a seminal paper, [HNWW00], an alternative upper bound for the smallest achievable star-discrepancy of the form
$D_{N}^{*}(P)\leq c\sqrt{\frac{d}{N}}$ (2)
@@
for all $d,N\in\mathbb{N}$ and $\varepsilon\in[0,1)$. Currently the best known value for the constant is $c=2.4968$ according to [GPW21]. This observation can also be used to check whether a given point set or sequence is good in the sense that its empirically observed star-discrepancy is close to or even smaller than the upper bound. As a special example for this idea it was shown in [GGPP21] that the sequence generated by a secure bit generator is good up to dimension at least $d=15$, because its discrepancy is smaller than $\sqrt{\frac{d}{N}}$ even for relatively big $N$.
@@
$(x_{n}):=(\varphi_{p_{1},\pi}(f_{1}(n)),\ldots,\varphi_{p_{d},\pi}(f_{d}(n)))$
@@
Thus, all scrambled Halton subsequences of the form $f(n)=(a_{1}\cdot n,a_{2}\cdot n,\ldots a_{d}\cdot n)$ with $\gcd(a_{i},p_{i})=1$ for $i=1,\ldots,d$ are low-discrepancy sequences. We call these $a_{i}$ shifts.
@@
Note that for the original Halton sequence, the pre-factor in front of $\log(N)^{d}/N$ is known to be smaller by $2^{2d}$, if we look at the star-discrepancy instead of the discrepancy, see e.g. [KN74].
Also in the multi-dimensional setting, the permutations $\pi_{i}$ and the shifts $a_{i}$ may be chosen in order to minimize the star-discrepancy. Our numerical results in this setting are promising and we obtain values for the discrepancy which are at least comparable to other recent methods for finding sets with small discrepancy, see [DR13], [CVJ+23]. As we want to analyze this case extensively but not overload the introduction, we postpone the detailed discussion to Section 2.
@@
Since finding the (empiric) small values for the (inverse) star-discrepancy requires an extensive search, it can only cover some cases $N,d\in\mathbb{N}$. As a theoretical application of Halton (sub-)sequences, they can be used to improve the value of $c$ in the theoretical bound (2). In fact, numerical calculations for finitely many $N$ can cover those cases for which the application of bounds like in Theorem 2 is not sufficient.
@@
Let $d,N\in\mathbb{N}$. Then there exists a point set $P\subset[0,1]^{d}$ such that
$D_{N}^{*}(P)\leq 2.4631837\sqrt{\frac{d}{N}}.$
The general structure of the proof for Theorem 4 is in parallel to [GPW21]. However, we add some additional ideas to improve the bound: we use Halton sequences and improved bounds on their discrepancy from [Ata04] to address the case $d\leq 4$ while only the case $d=1$ had been addressed separately before. Moreover, we were able to slightly sharpen the arguments to derive the bounds. Finally, we use a very recent result on bracketing numbers from [Gne24]. We will discuss the details in Section 3.
It is natural to ask why our approach is only applied to the case $d\leq d_{0}=4$. From a theoretic viewpoint, there is no reason for this choice and we would expect that (scrambled) Halton (sub-)sequences could probably be used for any finite $d_{0}$: given $d_{1}\leq d_{0}$, the bound from [Ata04] can be applied to prove an analogue of Theorem 4 for all but finitely many $N\in\mathbb{N}$. However, these finitely many exceptions need to be checked on a computer.
In [GSW09] it was proven that calculating the (star-)discrepancy is an NP-hard problem. Indeed, all known algorithms for calculating the star-discrepancy have exponential run time. The currently best known one was introduced in [DEM96] and is also called DEM algorithm named after its inventors Dobkin, Epstein and Mitchell. Its runtime is of the order $O(N^{1+d/2})$, where $N$ is the cardinality of the set, compare [CVJ+23]. For our calculations in higher dimensions, see Section 2, we use a more recent implementation of the DEM algorithm based on earlier work of Magnus Wahlström publicly available under [CVdN+23]. We do not describe its details here but refer to the description of the DEM algorithm in [DGW14].
@@
The remainder of the paper is organized as follows: In Section 2, we discuss numerical results for the application of Theorem 2 in a multi-dimensional setting. We will show for many different combinations of $d,N\in\mathbb{N}$ that we obtain sets with a star-discrepancy which is at most as big as the currently best known ones obtained by alternative methods. Afterwards, we will give proofs of our theoretical results in Section 3. The proofs of Theorems 1 and Theorem 2 rely on p-adic discrepancy theory as discussed in e.g. [Wei24]. Furthermore, Theorem 4 is obtained by using Halton (sub-)sequences in small dimensions and an improved bound on bracketing numbers proven in [Gne24]. Finally, we note the explicit combinations of primes, shifts and permutations by which we obtain our optimal numerical results in the Appendix 4. This puts the reader into the position to reproduce our calculations.
@@ Acknowledgments.
The author would like to thank Francois Clément for discussions on the content of the paper and for providing the link to the C-code for performing the calculations of the DEM algorithm. Moreover, he is grateful to Michael Gnewuch for his comments on a preliminary version of this paper and especially on the proof of Theorem 4 .

### v2

@@ Abstract
Braaten and Weller discovered that the star-discrepancy of Halton sequences can be strongly reduced by scrambling them. In this paper, we apply a similar approach to those subsequences of Halton sequences which can be identified to have low-discrepancy by results from $p$-adic discrepancy theory. For given finite $N$, it turns out that the star-discrepancy of these sequences is surprisingly low. By that known empiric bounds for the inverse star-discrepancy can be improved. Furthermore, we establish the existence of $N$-point sets in dimension $d$ whose star-discrepancy satisfies $\leq 2.4631832\sqrt{\frac{d}{N}}$, where the constant improves upon all previously known bounds.
@@ 1 Introduction
i.e. the intervals in the supremum are not necessarily anchored at 00. It relates to the star-discrepancy via $D_{N}^{*}(y_{n})\leq D_{N}(y_{n})\leq 2D_{N}^{*}(y_{n})$. Thus, the two notions necessarily possess the same asymptotic behavior up to a factor.
@@
Recall that for an integer $b\geq 2$ the $b$-ary representation of $n\in{\mathbb{N}}$ is $n=\sum_{j=0}^{\infty}e_{j}b^{j}$ with $0\leq e_{j}=e_{j}(n)<b$. Based on the radical-inverse function, the van der Corput sequence in base $b$ is defined by $\varphi_{b}(n)=\sum_{j=0}^{\infty}e_{j}b^{-j-1}$ for all $n\in{\mathbb{N}}_{0}$. It is well-known, see e.g. [Nie92], that
$D_{N}^{*}(\varphi_{b}(n))\leq\frac{1}{N}+\frac{b+1}{2N}+\frac{b-1}{2\log(b)}\frac{\log(N)}{N}$
This bound should be compared to (and is for small $b$ not too far off from) the record holder for the best known star-discrepancy as constructed in [Ost09] which satisfies
@@
In fact, this record holder is a generalized van der Corput sequence in base $b=60$. In other words, van der Corput, their subsequences and generalizations may be regarded as prime candidates when looking for sequences with a particularly small star-discrepancy.
An easier-to-describe method than in [Ost09] how to (empirically) further reduce the star-discrepancy of van der Corput sequences was analyzed in detail in [BW79] and was based on an idea from [Fau78]: for fixed $b\in\mathbb{N}$ choose an arbitrary permutation $\pi$ of $\{0,\ldots,b-1\}$ and define the scrambled van der Corput sequence by
@@
It is not difficult to prove the same bound for the star-discrepancy as for the standard van der Corput sequence but the additional parameter allows for (empiric) reduction of the star-discrepancy. In [BW79], concrete choices for $\pi$ were suggested for all prime bases $p\leq 53$.
Actually, the result in [Ost09] relies on the (rather complicated) exact formulae for the star-discrepancy of scrambled van der Corput sequences derived in [Fau81]. The article [Fau81] also provides asymptotic constants for the limit superior of both the extreme discrepancy and the star discrepancy. As another main theoretical result it was proven in [Fau92], that for every base $b$ there exists a permutation $\pi$ such that the constant in the inequality $D_{N}^{*}(\varphi_{b,\pi}(n))\leq c\frac{\log(N)}{N}$ is less than $1/\log(2)$.
The concept of scrambling can, of course, also be applied to subsequences of van der Corput sequences and we obtain the following result which may be regarded as a generalization of Theorem 1.2 in [Wei25] in combination with Theorem 3 in [Mei68] as will become clear from its proof. Recall that a polynomial $f:\mathbb{Z}\to\mathbb{Z}$ is called a permutation polynomial $\textrm{mod}\ n$ for $n\in\mathbb{N}$ if it is a bijection on $\mathbb{Z}/n\mathbb{Z}$ and thus a permutation of the elements in $\mathbb{Z}/n\mathbb{Z}$.
@@
Let $f(n)$ be a permutation polynomial $\bmod\ p^{2}$ for some prime number $p$, which means that $f(n)$ induces a bijection on $\mathbb{Z}/p^{2}\mathbb{Z}$, and let $\pi$ be a permutation of $\{0,\ldots,p-1\}$. Then the extreme discrepancy of the sequence $(x_{n}):=(\varphi_{p,\pi}(f(n)))$ satisfies
@@
We call such sequences scrambled van der Corput subsequences. The easiest examples are of the form $f(n)=an+b$ with $\gcd(a,p)=1$. We call $a$ a shift and denote such a sequence by $\varphi_{p,\pi,a}$. Other possible choices for $f(n)$, depending on $p$, may be found in [Wei25]. Although the theoretical bound does not guarantee a very small extreme or star-discrepancy, allowing the flexibility in choosing the parameters $\pi$ and $a$ in Theorem 1 reduces the extreme or star-discrepancy for given $N$ significantly as can be seen from Table 1 where our results are compared to the approach from [BW79]. The difference between the original van der Corput sequence and the scrambled one according to [BW79] is much larger (except for $p=2$), but moving to subsequences always but once reduced the star-discrepancy as well. In our experiments, the gain exceeded $10\%$ most of the time.
@@
Table 1: Comparison of star-discrepancies for van der Corput like sequences for $N=100$ (left) and $N=1,000$ (right). For $\varphi_{p,\pi,a}(N)$ the smallest found values are listed when using $500$ different shifts, $500$ random permutations per shift.
Finding sequences with a particularly small (star-)discrepancy is not only a relevant question in dimension $d=1$ but even of higher importance for $d>1$, where the star-discrepancy for a sequence $(x_{n})\subset[0,1]^{d}$ and the $d$-dimensional Lebesgue measure $\lambda_{d}(\cdot)$ is defined by
@@
where the supremum is taken over all $d$-dimensional intervals $B=[0,b_{1})\times\ldots\times[0,b_{d})$ with $0\leq b_{i}\leq 1$ for $i=1,\ldots,d$. It is widely conjectured that $D_{N}^{*}(x_{n})=O(\log(N)^{d}/N)$ is the best achievable order of convergence for the star-discrepancy but this conjecture has only been proven in the case $d=1$, see [Sch72]. Furthermore, we remind the reader that the star-discrepancy of sequences (infinitely many points) in dimension $d$ corresponds to that of point sets (finitely many points) in dimension $d+1$, see [Rot54]. The extreme discrepancy $D_{N}(x_{n})$ again allows arbitrary multi-dimensional half-open intervals in the supremum, i.e. without necessarily anchoring one base point at zero. As a generalization of the one-dimensional case, the inequality $D_{N}^{*}(x_{n})\leq D_{N}(x_{n})\leq 2^{d}D_{N}^{*}(x_{n})$ holds.
Sequences with a particularly small (star-)discrepancy are of great interest for high-dimensional integration tasks. Due to the Koksma-Hlawka inequality, see e.g. [KN74], the worst case error when approximating an integral by the average of some function evaluations depends linearly on the star-discrepancy of the evaluation points. This motivates why it makes sense to ask, what is the smallest star-discrepancy achievable for a given $N\in\mathbb{N}$ and to set
$D^{*}(N,d):=\inf\{D_{N}^{*}(P)\,:\,P\subset[0,1]^{d},\#P=N\}.$
Furthermore, define the inverse star-discrepancy by
@@
which is the minimum number of sample points that guarantees a star-discrepancy bound of at most $\varepsilon>0$. Note that even small reductions of $N$ are of practical importance if an expensive function $f$ is evaluated at these points to (numerically) calculate an integral. Alternatively, decreasing the size of $N$ can be motivated by the numerical stability of certain regression problems which depend linearly on the star-discrepancy, see [WN19]. Therefore, precise theoretical bounds and numerical estimates of $N^{*}(\varepsilon,d)$ are of great interest.
As the asymptotic bound $\log(N)^{d}/N$ for the star-discrepancy exponentially depends on the dimension, the enumerator is large in comparison to the denominator if $N$ is small. Therefore, the bound is hardly of any use in this setting. In the seminal paper, [HNWW00], an alternative upper bound for the smallest achievable star-discrepancy of the form
$D^{*}(N,d)\leq c\sqrt{\frac{d}{N}}$ (2)
@@
for all $d,N\in\mathbb{N}$ and $\varepsilon\in[0,1)$. Currently, the best known value for the constant is $c=2.4968$ according to [GPW21]. This observation can also be used to check whether a given point set or sequence is good in the sense that its empirically observed star-discrepancy is close to or even smaller than the upper bound. As a special example for this idea, it was shown in [GGPP21] that the sequence generated by a secure bit generator is good up to dimension at least $d=15$, because its star-discrepancy is smaller than $\sqrt{\frac{d}{N}}$ even for relatively large $N$.
@@
$(x_{n}):=(\varphi_{p_{1},\pi_{1}}(f_{1}(n)),\ldots,\varphi_{p_{d},\pi_{d}}(f_{d}(n)))$
@@
Thus, all scrambled Halton subsequences of the form $f(n)=(a_{1}\cdot n,a_{2}\cdot n,\ldots a_{d}\cdot n)$ with $\gcd(a_{i},p_{i})=1$ for $i=1,\ldots,d$ are low-discrepancy sequences. We also call these $a_{i}$ shifts.
@@
Note that for the original Halton sequence, the pre-factor in front of $\log(N)^{d}/N$ is known to be smaller by $2^{2d}$, if we look at the star-discrepancy instead of the extreme discrepancy, see e.g. [KN74].
Also in the multi-dimensional setting, the permutations $\pi_{i}$ and the shifts $a_{i}$ may be chosen in order to minimize the star-discrepancy. Our numerical results in this setting are promising and we obtain values for the star-discrepancy which are at least comparable to other recent methods for finding sets with small star-discrepancy, in particular[DR13], [CVN+23]. As we want to analyze this case extensively but not overload the introduction, we postpone the detailed discussion to Section 2.
@@
Since finding (empiric) small values for the (inverse) star-discrepancy requires an extensive search, it can only cover some cases $N,d\in\mathbb{N}$. As a theoretical application of scrambled Halton (sub-)sequences, they can be used to improve the value of $c$ in the theoretical bound (2). In fact, numerical calculations for finitely many $N$ can cover those cases for which an application of bounds like in Theorem 2 is not sufficient.
@@
For any $d,N\in\mathbb{N}$ we have
$D^{*}(N,d)\leq 2.4631832\sqrt{\frac{d}{N}}.$
The general structure of the proof for Theorem 4 is in parallel to [GPW21]. However, we add some additional ideas to improve the bound: First, we were able to slightly sharpen the arguments to derive the bounds. Second, we include a very recent result on bracketing numbers from [Gne24]. Last but not least, we employ Halton sequences along with improved bounds on their star-discrepancy from [Ata04] to address the case $d\leq 4$ whereas previously only the case $d=1$ had been treated separately. It turns out that the star-discrepancy of finitely many Halton sequence points must be computed individually for each dimension $d$ to complete the proof. Without utilizing the result from [Ata04] the number of necessary computations in dimension 4 would be prohibitively large. Even with this result, additional arguments are needed to further reduce the computational effort, making it feasible to carry out the calculations on a standard computer. This double reduction in computational complexity constitutes the main new contribution of the present article. We will discuss the details in Section 3, see in particular Remark 13.
It is natural to ask why our approach is only applied to the case $d\leq d_{0}=4$. From a theoretic viewpoint, there is no reason for this choice and we would expect that (scrambled) Halton (sub-)sequences could be used for any finite $d_{0}$: given $d_{1}\leq d_{0}$, the bound from [Ata04] can be applied to prove an analogue of Theorem 4 for all but finitely many $N\in\mathbb{N}$. However, these finitely many exceptions need to be checked on a computer, which is a very demanding task, especially in higher dimensions.
In [GSW09] it was proven that calculating the (star-)discrepancy is an NP-hard problem. Indeed, all known algorithms for calculating the star-discrepancy have exponential run time. The currently best known one was introduced in [DEM96] and is also called DEM algorithm named after its inventors Dobkin, Epstein and Mitchell. Its runtime is of the order $O(N^{1+d/2})$, where $N$ is the cardinality of the set, compare [CVN+23]. For our calculations in higher dimensions, see Section 2, we use a more recent implementation of the DEM algorithm based on earlier work of Magnus Wahlström publicly available under [CVdN+23]. We do not describe its details here but refer to the description of the DEM algorithm in [DGW14].
@@
The remainder of the paper is organized as follows: In Section 2, we discuss numerical results in a multi-dimensional setting, which are based on Theorem 2. We will show for many different combinations of $d,N\in\mathbb{N}$ that we obtain sets with a star-discrepancy which is at most as big as the currently best known ones obtained by alternative methods. Afterwards, we will give proofs of our theoretical results in Section 3. The proofs of Theorems 1 and Theorem 2 rely on $p$-adic discrepancy theory as discussed in e.g. [Wei25]. Furthermore, Theorem 4 is obtained by using Halton (sub-)sequences in small dimensions and an improved bound on bracketing numbers proven in [Gne24]. Finally, we collect the explicit combinations of primes, shifts and permutations by which we obtain our optimal numerical results in the Appendix 4. This puts the reader into the position to reproduce our calculations.
@@ Acknowledgments.
The author would like to thank Francois Clément for discussions on the content of the paper and for providing the link to the C-code used to perform the calculations of the DEM algorithm. Moreover, he is grateful to Michael Gnewuch for his comments on a preliminary version of this paper and especially on the proof of Theorem 4.